In a certain senior class, 72 percent of the male students and 80 percent of the female students have applied to college. What fraction of the students in the senior class are male?
(a) There are 840 students in the senior class
(b) 75 percent of the students in the senior class have applied to college

Found 2 solutions by ewatrrr, checkley77:
Answer by ewatrrr(24785)


Hi,
There are 840 students in the senior class
75 percent of the students in the senior class have applied to college
Let x represent the number of females. then (840 -x ) represents the number of males.
question states
.72(840-x) + .80x = .75*840
solving for x
.08x = .03*840
x = 315 of the senior class are females
840-315 = 525 are males
checking our answer
.72*525 + .80*315 = 378 + 252 = .75*840 = 630

Answer by checkley77(12844)

.72x+.80(840-x)=.75*840
.72x+672-.80x=630
-.08x=630-672
-.08x=42
x=42/.08
x=525 are boys.
525/840=5/8 are boys.
---------------------------------
840-525=315 are girls.
Proof:
.72*525+.80*315=.75*840
378+252=630
630=630